QUANTUM ANTIGRATIVITION OF VAPOUR-AIR JET

\underline{KHOLMURAD KHASANOV}, Samarkand University, 15, University Boulevard, Samarkand, Uzbekistan, 703004 (e-mail:hanvar@samarkand.uz).

We invented the extra-sound device (ESD) which allows to receive the stable extra-sound gas jet with regular interacting structure-shock waves in flooded space. In present paper results of new experiments: beginning from some initial value of pressure in the vicinity of the radius r ∼ 10m of the water-vapour jet in air there arises the phenomenon of quantum antigravitation generated by regular waves. In this vicinity of vapour-air jet there defined the parameters of electromagnetic wave

ν = 10^{9} − 10^{16}Hz  λ : 4 ∗ 10^{−2} − 3 ∗ 10^{−8}m;  (1)

also, were defined: voltage from 1kV to 50kV, electric field strength ∼ 10^{4} \frac{V}{m}, strength in vapour-air jet ∼ 10^{−6}/10^{−3}A, power of quant radiation at electrons density 10^{15}m^{−3} is 10^{5}1/c, Lengmure’s frequency, wavelength of which corresponds to the experimmental data. We have noted that the character of quants radiation on distance 2-3 radius defined above has antigravitational direction and the potential in the 20-30m distance from jet is constant. Radiation is directed from Earth’s surface upwards.